OpenStudy (anonymous):

Plz help, got like 20 min left!!! What is the axis of symmetry of the graph of y = x^2- 3x + 9?

OpenStudy (anonymous):

x=3 axis of symmetry is x coordinate of vertex

OpenStudy (anonymous):

x=3 axis of symmetry is x coordinate of vertex. lml yea what he said ^^

OpenStudy (anonymous):

Thanks yall are lifesavers

OpenStudy (anonymous):

Nope, it's not 3. It's x = -b/(2a) = -(-3) / (2*1) = 3/2

OpenStudy (anonymous):

YES!!!!! i GOT A 100%!!!!!!!!!

OpenStudy (pixiedust1):

congratulations!! :D

OpenStudy (anonymous):

ON A UNIT TEST!!!!!

OpenStudy (anonymous):

IN A CLASS THAT I AM TOTALLY FAILING!!!!

OpenStudy (pixiedust1):

Terrific! Those count big-time, right?

OpenStudy (anonymous):

yeah, it brought my grade up 5 points

OpenStudy (pixiedust1):

Yeah! Are you passing now?

OpenStudy (anonymous):

not yet.....need like 3 more points

OpenStudy (pixiedust1):

Getting close lol! If you need any more help, just let me know :)

OpenStudy (anonymous):

okay thank you ill message you if i do.